OpenStudy (anonymous):

Two similar prisms have heights 4 cm and 10 cm. What is the ratio of their surface areas? A) 2:5 B) 4:25 C) 4:10 D) 8:125 Thanks!

OpenStudy (anonymous):

What does it mean if they're 'similar' ?

OpenStudy (anonymous):

Would this be A?

OpenStudy (anonymous):

Their heights reduced?

OpenStudy (anonymous):

Well that's a good place to start. However, what does it mean that they are 'similar' ?

OpenStudy (anonymous):

It means that they are comparable...

OpenStudy (anonymous):

Specifically it means that the ratio of their dimensions will be the same right?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

aha so c?

OpenStudy (anonymous):

If the heights have a ratio of 2:5, the dimensions of their bases will also have a ratio of 2:5.

OpenStudy (anonymous):

I'm lost... obviously

OpenStudy (anonymous):

Ok look. You know that the base of the prisim has some dimensions.. maybe it's a square, maybe it's a rectangle. Whatever it is, the dimensions (width and height) will have a 2:5 ratio with the base of the larger prism.

OpenStudy (anonymous):

So if the area of the base is \(L \times W\) then the ratio of the area will be : \[2L \times 2W \over 5L \times 5W\] So it's going to have a 4:25 ratio in area.

OpenStudy (anonymous):

Because it has a 2:5 ratio in dimension.

OpenStudy (anonymous):

okay, so just times the dimensions by themselves? Or was that just with this problem? Thank you!

OpenStudy (anonymous):

If you want I can try for a more detailed explanation that might make it easier to understand.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

That was my initial guess, but polak worked it out and said 4.25?

OpenStudy (anonymous):

Ok, look (since this will be helpful in the next problem): Let \(W_1\) be the width of the smaller prisms base. Let \(L_1\) be the length of the smaller prisms base. Let \(H_1\) be the height of the smaller prism. Let \(W_2\) be the width of the larger prisms base. Let \(L_2\) be the length of the larger prisms base. Let \(H_2\) be the height of the larger prism. So we have that the two are similar and that: \[H_1 = \frac{2}{5}H_2\] Therefore (because they are similar) \[L_1 = \frac{2}{5}L_2\]\[W_1 = \frac{2}{5}W_2\] Therefore the area of the bases will be \[A_1 = C(L_1 \times W_1)\] Where C is some constant pertaining to the shape of the base (1/2 for triangles, etc) Therefore: \[A_2 = C(L_2\times W_2)\] And their ratios are \[\frac{A_1}{A_2} = {C(L_1 \times W_1) \over C(L_2\times W_2)}\]\[= {L_1 \times W_1 \over L_2\times W_2}\]\[={\frac{2}{5}L_2 \times \frac{2}{5}W_2 \over L_2\times W_2}\]\[=\frac{2}{5}\times\frac{2}{5} \times {L_2 \times W_2 \over L_2\times W_2} = {4 \over 25}\]

OpenStudy (anonymous):

Thanks so much for helping me! I wish I could understand all this though, lol :(

OpenStudy (anonymous):

Which part is confusing you?

OpenStudy (anonymous):

I have trouble comprehending just how exactly I would do those steps on paper, or with my calculator...

OpenStudy (anonymous):

No calculator required. You'd do it on paper just the way I did. The question is, do you understand what I did?

OpenStudy (anonymous):

If not, which part doesn't make sense?

OpenStudy (anonymous):

We can go through it more slowly.